Is it possible to have a triangle with the following sides?
​3 cm, 6 cm, 7 cm

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Clearly, we have:
(3 + 6) = 9 > 7
(6 + 7) = 13 > 3
(7 + 3) = 10 > 6
Thus, the sum of any two of these numbers is greater than the third.
Hence, it is possible to draw a triangle whose sides are 3 cm, 6 cm and 7 cm.

Thanks